Calculate calorie to nitrogen ratio of a TPN solution consisting of 1500 ml of 50% dextrose, 10% fat, and 8.5% amino acid?

A. 216.2

B. 11.6

C. 191.2

D. 25.0

E. 30.6
Answer #1

C.191.2

Calculation of calories:

50% dextrose - 50g/100ml = X/1500ml

                                      X    = 75g

Calories in 50% dextrose = 75g * 3.4 kcal/g = 2550kcal

10% fat - 10g/100ml = X/1500ml

                                      X     = 15g

Calories in 10% fat = 15g * 9 kcal/g = 1350 kcal

Total nonprotein calories = 2550kcal+1350kcal = 3900kcal

Calculation of grams of nitrogen in protein(1g N = 6.25g):

8.5% protien - 8.5g/100ml = X/1500ml

                                     X    = 127.5g

Grams of nitrogen in 8.5% protein = 127.5g/6.25g = 20.4

Calculation of calorie to nitrogen ratio:

Divide the nonprotein calories by grams of nitrogen

Calorie to nitrogen ration = 3900/20.4 =191.2